1. If 75ml of AgNO3 solution reacts with enough Cu to produce 0.25g of Ag by single displacement, what is the molarity of the initial AgNO3 solution ?for the reaction AgNO3+Cu=Cu(NO2)2+Ag
1
Expert's answer
2021-02-26T05:58:53-0500

Solution:

The balanced chemical equation:

2AgNO3 + Cu → Cu(NO3)2 + 2Ag

According to the equation above: n(AgNO3) = n(Ag)


Moles of Ag = Mass of Ag / Molar mass of Ag

The molar mass of Ag is 107.8682 g mol-1.

Hence,

n(Ag) = 0.25 g / 107.8682 g mol-1 = 0.00232 mol


n(AgNO3) = n(Ag) = 0.00232 mol


Molarity of AgNO3 = Moles of AgNO3 / Volume of AgNO3 solution

Molarity of AgNO3 = 0.00232 mol / 0.075 L = 0.0309 mol/L = 0.0309 M

Molarity of AgNO3 = 0.0309 M


OR:

(0.25 g Ag / 75×10-3 L soln.)×(1 mol Ag / 107.8682 g Ag)×(2 mol AgNO3 / 2 mol Ag) = 0.0309 M AgNO3


Answer: The molarity of the initial AgNO3 solution is 0.0309 M.